# How to be an orthodox quantum mechanic

### v1

@@ 3 Sample of studied works
To perform this study we examine the $33$ popular textbooks listed in Table 1. These range in date of first publication from 1930 to 2019. This large timespan is of utility, as it allows us to determine whether there is any evolution within the consensus view of quantum mechanics. Additionally, we consider subdivision of these works into graduate ($N=14$) and undergraduate ($N=19$), as determined from the statements of the target audience in the work itself. These works were chosen to encompass the Wikipedia list on ‘Notable quantum mechanics textbooks’ [^2] https://en.wikipedia.org/wiki/List_of_textbooks_on_classical_mechanics_and_quantum_mechanics ($N=17$) as well as to include as large a sample as feasible. To detect variation over time we also consider a sub-sample of works first published after 1980 ($N=24$). Note that Omnés work is included, even though it is a textbook that advocates against the orthodoxy in favour of “consistent histories”, as it provides its own description of the orthodox positions. In the case of [13], this thoughtfully discusses some foundational issues without expressing a direct opinion, here we substitute what the author characterises as the ‘conventional” approach. We will also make reference to a set of 4 works whom we do not include in any of the samples: [4, 14, 15, 12]. The first is excluded as it is a discussion of interpretations but will provide a useful point of comparison. The second two are analysed out of curiosity but are not targeted squarely as part of a physics syllabus. The final one will be analysed and found to be entirely outside the orthodoxy on almost every point of contention (so it is treated as an outlier).
@@
Quantum physics of atoms, molecules, R. Eisberg & R. Resnick 1974 [37] U Y
solids, nuclei and particles
@@ 4 Analysis from example texts
To illustrate the analysis process we will consider several texts as examples. These are chosen because they all share a common case of having to infer agreement with at least one proposition (usually “measurement creates observed values”). This will allow the reader to determine that we are carefully evidencing degrees of assent or dissent in each case. Note that the sections below leave out propositions that are unaddressed by the given text.
@@ 4.1 Binney & Skinner
We begin with The physics of quantum mechanics by Binney & Skinner [18]. For “measurement disturbs systems” we find initial evidence on page 3: “Moreover, the act of measurement inevitably disturbs the atom, and leaves it in a different state from the one it was in before we made the measurement” (note that this is not in the context of the physics of a particular type of measurement). This is reinforced on page 130 in a discussion of the collapse hypothesis: “The Copenhagen interpretation does, however, contain a crucial insight into measurement by stressing that any measurement physically disturbs the system, so the system’s state after a measurement has been made is different from what it was earlier”. For “uncertainty linked to disturbance” we find evidence for endorsement on page 3 (“We rarely know the exact state that our measuring instrument is in before we bring it into contact with the system we have measured, so the result of the measurement of the atom would be uncertain even if we knew the precise state that the atom was in before we measured it, which of course we do not”), page 28 (a discussion of the uncertainty principle in terms of measurement uncertainty) which is subsequently reinforced on pages 130 and 131. For single particle uncertainty we find on page 28 “the position of a particle that has well defined momentum is maximally uncertain” as evidence for adherence to this position. For “particles are wave-like” we find a discussion opposing this on page 7, particularly “we cannot consistently infer from this correspondence that particles are manifestations of waves”. For “collapse”, the work opposes this idea explicitly stating collapse is not physical on page 15 and that the hypothesis is false on page 130. For “particle-wave duality” the authors never mention the concept explicitly but already oppose “single particles are wave-like” outside of quantum field theory contexts (page 7). For “non-ensemble $\psi$” we find on page 14 “in quantum mechanics an individual photon has an amplitude to be linearly polarised in a any chosen direction and an amplitude to be circularly polarised in a given sense”, so states correspond to individual system realisations. This is reinforced on page 71 where they state “In quantum mechanics, complete information about any system is contained in its ket $\left|\psi\right>$”. For “probability from uncertainty” the authors dissent, as we have seen that they attribute probability to the uncertain state of the measuring apparatus and its subsequent disturbing of the system. For “measurement creates observed values” we infer assent, their view seems to be that systems are disturbed by measurement, changing the values found rather than creating them altogether. We have already seen that the authors view the state as ontologically complete, an “extended” theory is possible, but it merely accounts for the actual physics of measurement (page 130 and 131).
@@ 4.2 Tamvakis
Now we consider Basic quantum mechanics by Tamvakis [16]. On page 86 it is outlined, as part of the axioms of quantum mechanics, that the state changes during measurement indicating support for “measurement disturbs systems”. On page 5 uncertainty is described in terms of single particle measurement precision or “if we make a measurement on any object and determine its position along a given direction with uncertainty $\Delta x$, its momentum along this direction will necessarily be known with uncertainty no less than $\frac{h}{4\pi\Delta x}$”, an indication of “single particle uncertainty”. On page 6 the de Broglie relation is discussed as the wave aspect of a single particle, implying individual particles are wave-like. Pages 86 and 87 discuss a non-unitary evolution during measurement and call this collapse, this re-iterated on page 442. On page 5: “the particle aspects of a system are complementary to its wave aspects” directly supports complementarity. On page 7 it is stated that “the wave function of the particle characterises the system fully at each point in time”. This indicates ontological completeness and a single realisation/particle association also supported by the previous quote. Page 5 describes the uncertainty principle in terms of knowledge of $x$ and $p_{x}$ “thus, viewed from a classical viewpoint, in a microscopic system, complete knowledge of $x$ and $p_{x}$ is impossible”, we infer this is indicating that probabilities result from incomplete knowledge implied by the uncertainty principle. For “measurement observed creates values” we note that it is stated on page 86 that a superposition represents the pre-measurement state, since this cannot have a definite property value we must conclude that definite property values are created, or at least changed, during measurement. This will satisfy an assent to “measurement observed creates values” as a pre-measurement state with definite values cannot be recovered as the author describes measurement as “acausal” on page 87.
@@ 4.3 Zettili
Next we consider Quantum Mechanics: Concepts and Applications by Zettili [25]. On page 24 “measurement interferes with the state of microscopic objects” and “it is the effects of interference by the equipment on the system which is the essence of quantum mechanics” on page 165 demonstrate support for measurement disturbance. On page 169 uncertainty is described as a result of measurement “perturbing” the system. “Single particle uncertainty” is affirmed on page 28 by “If the $x$ component of the momentum of a particle is measured with an uncertainty $\Delta p_{x}$, then its $x$ position cannot, at the same time, be measured more accurately than $\Delta x=\frac{\hbar}{2\Delta p_{x}}$”. Support for single particles being wave-like is given on page 18: “each material particle behaves as a group of waves”. Pages 24 and 25 indicate “we cannot follow a microscopic particle along its motion, nor can we determine its path”, while 51 discusses wave packet spreading only meaning the “position cannot be known exactly”. These are taken to indicate trajectories exist but are unknown. On page 158 collapse is included in the postulates of quantum mechanics. On page 26 “Microscopic systems, therefore, are neither pure particle nor pure waves, they are both” as well as direct invocations on the same page, suffices for duality and complementarity. For ontological completeness, page 158 asserts “the state contains all the needed information about the system”. On page 30 we find “In quantum mechanics the state (or one of the states) of a particle is described by a wave function $\psi(r,t)$ corresponding to the de Broglie wave of this particle; so $\psi(r,t)$ describes the wave properties of a particle” indicates that the state applies to single instances. We combine previous statements from 165 and 169 to conclude that probability is introduced by measurement disturbance (uncertainty). Based on the axioms from 158 we infer that measurement creates observed values, as pre-measurement states can be superpositions and apply to individual particles/realisations.
@@ 4.4 Mcintyre, Manogue, & Tate
We now turn to Quantum mechanics by McIntyre et al [20]. On page 8 we find that measurements disturbing systems is a “key feature” of quantum mechanics. On page 59, in the context of the uncertainty principle, it is stated that “the lack of ability to measure all spin components simultaneously implies that the spin does not really point in a given direction”. This indicates the uncertainty principle is being applied to individual particles. Combining this with similar discussion on 208 (about position and momentum) implies a link between uncertainty and disturbance. A confusing discussion on 207 about the double-slit seems to suggest both that a wave picture is just for finding particle probabilities but also that the waves have to actually interfere (passing through two slits). The same discussion refers to Feynman and Cohen-Tannoudji which seems to indicate the ultimate outcome is individual particles are wave-like too. Page 192 indicates we cannot know a particles trajectory but does not deny its existence. Page 48 includes collapse in the postulates. Wave-particle duality is explicitly invoked as a correct description of quantum systems on pages 121 and 207. Complementarity in terms of position and momentum is discussed on pages 192 and broadened to particle and wave properties in 207 and 208. “The state of quantum mechanical system, including all the information you can know about it, is represented mathematically by a normalised ket $\left|\psi\right>$” covers ontological completeness. Discussion of Schrödinger’s cat on 167 and 168 is illuminating in terms of the role of ensembles. Here it is asserted that only decoherence prevents superpositions at classical scales and interference is used to ascribe physical signficance to superposition, beyond mere statistics. Thus, we see that the authors take superpositions to describe the state of individual realisations. Acceptance of real superpositions combined with the stated quantum postulates requires definite values are created by measurement.
@@ 4.5 Konishi & Paffuti
Finally we review Quantum mechanics: a new introduction by Konishi & Paffuti [21]. On page 29 the authors support the notion that all measurement causes disturbance by stating that “the disturbance caused by the process of measurement on the microscopic system under examination cannot be controlled beyond a certain limit.” The unqualified illustration of the uncertainty principle via “Heisenberg microscopes” on page 28 indicates the authors link uncertainty to measurement disturbance. On page 26 we find “the classical concept of a particle trajectory no longer makes sense. An electron cannot simultaneously have a definite position and momentum.” This demonstrates both that the uncertainty principle is applied to individual particles and that the authors reject trajectories for quantum objects. This comes with a caveat on page 29 that uncertainty principle is not itself fundamental, but derives from the “dynamical properties of the physical processes involved in the measurement.” On page 6 the authors indicate that single particles are wave-like by asserting that wave-particle duality “concerns the properties of an individual electron” and not a population. This, combined with page 28 highlighting the importance of wave-particle duality to uncertainty relations, supports a more general duality thesis too. On pages 31 and 32 we find that the authors believe state reduction is “necessary for the theory to make sense at all.” This is furthered on 518 and 519 where they argue collapse is experimentally confirmed. On page 25 the authors describe a superposed polarisation state as “the state of a single photon”. Additionally, on page 23 the authors ascribe “physical properties” to superposition states that are “somewhere between” those of the constituents. These indicate that quantum states do not apply to ensembles but to single realisations of systems. On page 29 the authors agree that probability derives from the uncertainty principle “the probabilistic nature of quantum mechanical pediction for the results of measurement is an inevitable consequence of Heisenberg’s uncertainty principle.” Although the discussion on 518 and 519 is ambiguous as to whether values are being revealed or created (the authors use “indefinite” properties of a superposition here to mean “unpredictable”), we must infer that they are at least being altered by measurement to comport with previous assertions. On page 21 the authors state that “knowledge of the wavefunction amounts to complete knowledge of the state” which, combined with prior assertions fully supports ontological completeness.
@@ 5 Results and discussion
In Table 2 we display all the data used for this study. Each work’s response to the propositions is characterised by a degree of agreement. This is kept as simple as possible with three principle values: $1$ indicates agreement, $-1$ dissent, and 00 as not mentioned. We expand this via either of ±$\pm 0.5$ to mean ‘inferred for logical consistency’. Inference of agreement/dissent is only made if the case is strong (in order to allow for a conservative analysis). We proceed to study the responses by observing the frequency of agreement. Note that this computation counts ±$\pm 0.5$ responses as half a count when determining frequencies. Our raw data keeps the page references separate for each proposition (available on request).
@@
In Figure 1, we see the agreement frequencies for each of the questions in turn. We display these in terms of 4 samples: the full sample (black), undergraduate books (yellow), graduate texts (blue), works on the ‘Notable books’ list from Wikipedia (red), and ‘modern’ works (gray). Full information on the degree of agreement, and the $p$ value computed from a binomial distribution are included in Tables 3 to 7. These $p$ values are drawn from the binomial distribution and represent the probability of the given degree of assent to arise by chance given the size of the sample. We choose a random binomial distribution because we consider a zero consensus result would be equal frequencies of assent ($1$) and non-assent (where we group 00 and $-1$ together). The grouping is done because no assertion in the sample has a consensus answer of 00 or $-1$. Thus, we take a high $1$ frequency as agreement with this assertion is “in the orthodoxy”. We also include a $\sigma$ value, the number of standard deviations of the binomial distribution that the assent frequency corresponds to.

### v2

@@ 3 Sample of studied works
To perform this study we examine the $42$ popular textbooks listed in Table 1. These range in date of first publication from 1930 to 2025. This large time span is of utility, as it allows us to determine whether there is any evolution within the consensus view of quantum mechanics. Additionally, we consider subdivision of these works into graduate ($N=17$) and undergraduate ($N=25$), as determined from the statements of the target audience in the work itself. These works were chosen to encompass the Wikipedia list on ‘Notable quantum mechanics textbooks’ [^2] https://en.wikipedia.org/wiki/List_of_textbooks_on_classical_mechanics_and_quantum_mechanics ($N=18$) as well as to include as large a sample as feasible. To detect variation over time we also consider three subsamples: works pre-1980 ($N=15$), post-1980 ($N=27$), and post-2000 ($N=17$). Note that Omnés work is included, even though it is a textbook that advocates against the orthodoxy in favour of “consistent histories”, as it provides its own description of the orthodox positions which we use for our study. In the case of [15], this thoughtfully discusses some foundational issues without expressing a direct opinion, here we substitute what the author characterises as the “conventional” approach. Note that page references for [16] are in reference to the online lecture notes on which it is based [^3] https://www.damtp.cam.ac.uk/user/tong/books/quantum.html, due to the recency of the text. We will also make reference to a set of 4 works whom we do not include in any of the samples: [6, 17, 18, 14]. The first is excluded as it is a discussion of interpretations, but will provide a useful point of comparison. The second two are analysed out of curiosity, but are not targeted squarely as part of a physics syllabus. The final one will be analysed and found to be entirely outside the orthodoxy on almost every point of contention (so it is treated as an outlier).
@@
Quantum mechanics for scientists and engineers D. Miller 2008 [27] U N
Quantum mechanics: a textbook for undergraduates M. Jain 2007 [28] U N
@@
Quantum physics of atoms, molecules, solids, nuclei and particles R. Eisberg & R. Resnick 1974 [46] U Y
@@ 4 Analysis from example texts
To illustrate the analysis process we will consider several texts as examples. These are chosen because they all share a common case of having to infer agreement with at least one proposition (usually “measurement creates observed values”). This allows the reader to determine that we are carefully evidencing degrees of assent or dissent in each case. Note that the sections below leave out propositions that are unaddressed by the given text.
@@ 4.1 Binney & Skinner
We begin with The physics of quantum mechanics by Binney & Skinner [22]. For “measurement disturbs systems” we find initial evidence on page 3: “Moreover, the act of measurement inevitably disturbs the atom, and leaves it in a different state from the one it was in before we made the measurement” (note that this is not in the context of the physics of a particular type of measurement). This is reinforced on page 130 in a discussion of the collapse hypothesis: “The Copenhagen interpretation does, however, contain a crucial insight into measurement by stressing that any measurement physically disturbs the system, so the system’s state after a measurement has been made is different from what it was earlier”. For “uncertainty linked to disturbance” we find evidence for endorsement on page 3 (“We rarely know the exact state that our measuring instrument is in before we bring it into contact with the system we have measured, so the result of the measurement of the atom would be uncertain even if we knew the precise state that the atom was in before we measured it, which of course we do not”), page 28 (a discussion of the uncertainty principle in terms of measurement uncertainty) which is subsequently reinforced on pages 130 and 131. For single particle uncertainty we find on page 28 “the position of a particle that has well-defined momentum is maximally uncertain” as evidence for adherence to this position. For “particles are wave-like” we find a discussion opposing this on page 7, particularly “we cannot consistently infer from this correspondence that particles are manifestations of waves”. For “collapse”, the work opposes this idea, explicitly stating collapse is not physical on page 15 and that the hypothesis is false on page 130. For “particle-wave duality” the authors never mention the concept explicitly but already oppose “single particles are wave-like” outside of quantum field theory contexts (page 7). For “non-ensemble $\psi$” we find on page 14 “in quantum mechanics an individual photon has an amplitude to be linearly polarised in any chosen direction and an amplitude to be circularly polarised in a given sense”, so states correspond to individual system realisations. This is reinforced on page 71 where they state “In quantum mechanics, complete information about any system is contained in its ket $\left|\psi\right$>”. For “probability from uncertainty” the authors dissent, as we have seen that they attribute probability to the uncertain state of the measuring apparatus and its subsequent disturbing of the system. For “measurement creates observed values” we infer assent, their view seems to be that systems are disturbed by measurement, changing the values found rather than creating them altogether. We have already seen that the authors view the state as ontologically complete, an “extended” theory is possible, but it merely accounts for the actual physics of measurement (page 130 and 131).
@@ 4.2 Tamvakis
Now we consider Basic quantum mechanics by Tamvakis [19]. On page 86 it is outlined, as part of the axioms of quantum mechanics, that the state changes during measurement indicating support for “measurement disturbs systems”. On page 5 uncertainty is described in terms of single particle measurement precision or “if we make a measurement on any object and determine its position along a given direction with uncertainty $\Delta x$, its momentum along this direction will necessarily be known with uncertainty no less than $\frac{h}{4\pi\Delta x}$”, an indication of “single particle uncertainty”. On page 6, the de Broglie relation is discussed as the wave aspect of a single particle, implying individual particles are wave-like. Pages 86 and 87 discuss a non-unitary evolution during measurement and call this collapse, this re-iterated on page 442. On page 5: “the particle aspects of a system are complementary to its wave aspects” directly supports complementarity. On page 7 it is stated that “the wave function of the particle characterises the system fully at each point in time”. This indicates ontological completeness and a single realisation/particle association also supported by the previous quote. Page 5 describes the uncertainty principle in terms of knowledge of $x$ and $p_{x}$ “thus, viewed from a classical viewpoint, in a microscopic system, complete knowledge of $x$ and $p_{x}$ is impossible”, we infer this is indicating that probabilities result from incomplete knowledge implied by the uncertainty principle. For “measurement creates observed values”, we note that it is stated on page 86 that a superposition represents the pre-measurement state, since this cannot have a definite property value we must conclude that definite property values are created, or at least changed, during measurement. This will satisfy an assent to “measurement observed creates values” as a pre-measurement state with definite values cannot be recovered, as the author describes measurement as “acausal” on page 87.
@@ 4.3 Zettili
Next we consider Quantum Mechanics: Concepts and Applications by Zettili [33]. On page 24, “measurement interferes with the state of microscopic objects” and “it is the effects of interference by the equipment on the system which is the essence of quantum mechanics” on page 165 demonstrate support for measurement disturbance. On page 169 uncertainty is described as a result of measurement “perturbing” the system. “Single particle uncertainty” is affirmed on page 28 by “If the $x$ component of the momentum of a particle is measured with an uncertainty $\Delta p_{x}$, then its $x$ position cannot, at the same time, be measured more accurately than $\Delta x=\frac{\hbar}{2\Delta p_{x}}$”. Support for single particles being wave-like is given on page 18: “each material particle behaves as a group of waves”. Pages 24 and 25 indicate “we cannot follow a microscopic particle along its motion, nor can we determine its path”, while 51 discusses wave packet spreading only meaning the “position cannot be known exactly”. These are taken to indicate trajectories exist but are unknown. On page 158 collapse is included in the postulates of quantum mechanics. On page 26 “Microscopic systems, therefore, are neither pure particle nor pure waves, they are both”, as well as direct invocations on the same page, suffices for duality and complementarity. For ontological completeness, page 158 asserts “the state contains all the needed information about the system”. On page 30 we find “In quantum mechanics the state (or one of the states) of a particle is described by a wave function $\psi(r,t)$ corresponding to the de Broglie wave of this particle; so $\psi(r,t)$ describes the wave properties of a particle” indicates that the state applies to single instances. We combine previous statements from 165 and 169 to conclude that probability is introduced by measurement disturbance (uncertainty). Based on the axioms from 158 we infer that measurement creates observed values, as pre-measurement states can be superpositions and apply to individual particles/realisations.
@@ 4.4 McIntyre, Manogue, & Tate
We now turn to Quantum mechanics by McIntyre, Manogue, and Tate [24]. On page 8 we find that measurements disturbing systems is a “key feature” of quantum mechanics. On page 59, in the context of the uncertainty principle, it is stated that “the lack of ability to measure all spin components simultaneously implies that the spin does not really point in a given direction”. This indicates the uncertainty principle is being applied to individual particles. Combining this with similar discussion on 208 (about position and momentum) implies a link between uncertainty and disturbance. A confusing discussion on 207 about the double-slit seems to suggest both that a wave picture is just for finding particle probabilities but also that the waves have to actually interfere (passing through two slits). The same discussion refers the reader to Feynman and Cohen-Tannoudji which seems to indicate the ultimate outcome is individual particles are wave-like too. Page 192 indicates we cannot know a particles trajectory but does not deny its existence. Page 48 includes collapse in the postulates. Wave-particle duality is explicitly invoked as a correct description of quantum systems on pages 121 and 207. Complementarity in terms of position and momentum is discussed on pages 192 and broadened to particle and wave properties in 207 and 208. “The state of quantum mechanical system, including all the information you can know about it, is represented mathematically by a normalised ket $\left|\psi\right$>” covers ontological completeness. Discussion of Schrödinger’s cat on 167 and 168 is illuminating in terms of the role of ensembles. Here it is asserted that only decoherence prevents superpositions at classical scales and interference is used to ascribe physical significance to superposition, beyond mere statistics. Thus, we see that the authors take superpositions to describe the state of individual realisations. Acceptance of real superpositions combined with the stated quantum postulates requires observed values are created by measurement.
@@ 4.5 Konishi & Paffuti
Finally we review Quantum mechanics: a new introduction by Konishi & Paffuti [26]. On page 29 the authors support the notion that all measurement causes disturbance by stating that “the disturbance caused by the process of measurement on the microscopic system under examination cannot be controlled beyond a certain limit.” The unqualified illustration of the uncertainty principle via “Heisenberg microscopes” on page 28 indicates the authors link uncertainty to measurement disturbance. On page 26 we find “the classical concept of a particle trajectory no longer makes sense. An electron cannot simultaneously have a definite position and momentum.” This demonstrates both that the uncertainty principle is applied to individual particles and that the authors reject trajectories for quantum objects. This comes with a caveat on page 29 that uncertainty principle is not itself fundamental, but derives from the “dynamical properties of the physical processes involved in the measurement.” On page 6 the authors indicate that single particles are wave-like by asserting that wave-particle duality “concerns the properties of an individual electron” and not a population. This, combined with page 28 highlighting the importance of wave-particle duality to uncertainty relations, supports a more general duality thesis too. On pages 31 and 32 we find that the authors believe state reduction is “necessary for the theory to make sense at all.” This is furthered on 518 and 519 where they argue collapse is experimentally confirmed. On page 25 the authors describe a superposed polarisation state as “the state of a single photon”. Additionally, on page 23 the authors ascribe “physical properties” to superposition states that are “somewhere between” those of the constituents. These indicate that quantum states do not apply to ensembles but to single realisations of systems. On page 29 the authors agree that probability derives from the uncertainty principle “the probabilistic nature of quantum mechanical prediction for the results of measurement is an inevitable consequence of Heisenberg’s uncertainty principle.” Although the discussion on 518 and 519 is ambiguous to whether values are being revealed or created (the authors use “indefinite” properties of a superposition here to mean “unpredictable”), we must infer that they are at least being altered by measurement to comport with previous assertions. On page 21 the authors state that “knowledge of the wave function amounts to complete knowledge of the state” which, combined with prior assertions fully supports ontological completeness.
@@ 5 Results and discussion
In Table 2 we display all the data used for this study. Each work’s response to the propositions is characterised by a degree of agreement. This is kept as simple as possible with three principle values: $1$ indicates agreement, 1-1 dissent, and 0 as not mentioned. We expand this via either of $\pm 0.5$ to mean ‘inferred for logical consistency’. Inference of agreement/dissent is only made if the case is strong (in order to allow for a conservative analysis). We proceed to study the responses by observing the frequency of agreement. Note that this computation counts $\pm 0.5$ responses as half a count when determining frequencies. Our raw data keeps the page references separate for each proposition (available on request).
@@
In Figure 1, we see the agreement frequencies for each of the questions in turn. We display these in terms of 4 samples: the full sample (black), undergraduate books (yellow), graduate texts (blue), works on the ‘Notable books’ list from Wikipedia (red), pre-1980 works (gray), post-1980 works (green), and post-2000 works (magenta). The only question substantially impacted by inference of response is “measurement” creates values. We supplement this plot with Fig. 2, which shows that rejection is not inferred to any significant degree. It would be ideal to include the significance of the consensus in each case. However, there is no obvious statistical prescription for doing so, as responses to the propositions are not generated by random processes in any sense. We note however, that many positions here achieve agreement frequencies above 80%. Given a sample of $42$ works, this is sign of substantial consensus.